1
___ - 3= 8
x(squared) _____
x

Answer by stanbon(75887)

1
___ - 3= 8
x(squared) _____
x
------------------------
[1/x^2] - 3 = [8/x]
Multiply thru by x^2 to get:
1 -3x^2 = 8x
Rearrange:
3x^2+8x-1 = 0
Use the Quadratic formula:
x = [-8 +- sqrt(64-4*3*-1)]/6
x = [-8 +- sqrt(76)]/6
x = -(4/3) + (1/3)sqrt19 or x = -(4/3) - (1/3)sqrt19
====================
Cheers,
Stan H.